On a 20-year mortgage due in 10 years the monthly mortgage payment include $2500 to fund a replacement reserve. What is the value of the replacement reserve at the end of three years assuming the reserve balace earns interest at 2% per annum, beginning balace is zero, and there are no withdrawls.

A) 92,757.64

B) 92,901.90

C) 92,829.73

D) 92,111.93
Answer #1

Monthly Replacement Reserve = $ 2500, Interest Rate = 2 % per annum or (2/12) = 0.167 % per month and Tenure = 3 years or (3x12) = 36 months

Accumulated Value of Reserve after 3 Years = 2500 x (1.00167)^(35) + 2500 x (1.00167)^(34) + ..........+ 2500 x (1.00167) + 2500 = 2500 x [{(1.00167)^(36)-1} / {1.00167 - 1}] = 2500 x 37.07229 = $ 92680.725 ~ $ 92757.64

Hence, the correct option is (A)